OpenStudy (anonymous):

Which of the following is an example of the associative property of multiplication? A. 7 • (3 + 5) = 21 + 35 B. 7 • (3 • 5) = (7 • 3) • 5 C. 7 • (3 • 5) = 7 • (5 • 3) D. 7 • = 1

OpenStudy (anonymous):

I don't get it

TheSmartOne (thesmartone):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 @Jason5679 Do you know what the associative property of multiplication is?

TheSmartOne (thesmartone):

Hint: \(\sf a \times (b \times c )= (a \times b) \times c\)